OpenStudy (anonymous):

Which is greater than seventeen twenty-fifths? zero point zero three five zero point three five zero point zero seven zero point seven

OpenStudy (anonymous):

zero point seven

OpenStudy (anonymous):

17/25 = 68/100 .7 = .70 = 70/100

OpenStudy (anonymous):

yes, but .7 =7/10 which is greater than 6/10(0.6) +8/100(.08) = 68/100?

OpenStudy (anonymous):

Sorry, read the question as less than, just realised my mistake, sorry